# AMEND: A Mixture of Experts Framework for Long-tailed Trajectory Prediction

### v1

@@ Abstract
Accurate prediction of pedestrians’ future motions is critical for intelligent driving systems. Developing models for this task requires rich datasets containing diverse sets of samples. However, the existing naturalistic trajectory prediction datasets are generally imbalanced in favor of simpler samples and lack challenging scenarios. Such a long-tail effect causes prediction models to underperform on the tail portion of the data distribution containing safety-critical scenarios. Previous methods tackle the long-tail problem using methods such as contrastive learning and class-conditioned hypernetworks. These approaches, however, are not modular and cannot be applied to many machine learning architectures. In this work, we propose a modular model-agnostic framework for trajectory prediction that leverages a specialized mixture of experts. In our approach, each expert is trained with a specialized skill with respect to a particular part of the data. To produce predictions, we utilise a router network that selects the best expert by generating relative confidence scores. We conduct experimentation on common pedestrian trajectory prediction datasets and show that besides achieving state-of-the-art performance, our method significantly performs better on long-tail scenarios. We further conduct ablation studies to highlight the contribution of different proposed components.
@@ I Introduction
Our contributions are as follows: We propose a novel mixture of experts framework for trajectory prediction. Our framework encourages the diversity of expert skill sets to mitigate the long-tailed distribution problem, while simultaneously avoids additional computational cost. We conduct empirical evaluations on common pedestrian trajectory benchmark datasets and show that our proposed approach achieves state-of-the-art performance. We conduct additional experiments, highlighting the advantage of our multi-expert method on predicting challenging scenarios. At the end, we perform ablation studies, showing the advantage of each proposed approach on the overall performance.
@@ IV-B Routing Samples to Experts
TABLE I: Quantitative evaluation on the UCY/ETH benchmark. Results are reported on $\text{minADE}_{20}$/$\text{minFDE}_{20}$. Bold numbers indicate the best performance for each metric, underline indicates the second best performance for each metric. For both metrics, lower value is better.
Method ETH Hotel Univ Zara1 Zara2 Avg
Traj++ EWTA [6] 0.38/0.65 0.10/0.17 0.18/0.34 0.14/0.27 0.10/0.19 0.18/0.32
MID [30] 0.39/0.66 0.13/0.22 0.22/0.45 0.17/0.30 0.13/0.27 0.21/0.38
Leapfrog [31] 0.39/0.58 0.11/0.17 0.26/0.43 0.18/0.26 0.13/0.22 0.21/0.33
NSP-SFM [32] 0.25/0.24 0.09/0.13 0.21/0.38 0.16/0.27 0.12/0.20 0.17/0.24
Y-Net [33] 0.28/0.33 0.10/0.14 0.24/0.41 0.17/0.27 0.13/0.22 0.18/0.27
SIC-MTP [34] 0.27/0.45 0.11/0.16 0.26/0.46 0.19/0.33 0.13/0.26 0.19/0.33
EQMotion [35] 0.40/0.61 0.12/0.18 0.23/0.43 0.18/0.32 0.13/0.23 0.21/0.35
AMEND (Ours) 0.29/0.42 0.08/0.12 0.13/0.21 0.12/0.20 0.09/0.14 0.14/0.22
@@
TABLE II: Quantitative evaluation on long-tail scenarios for the ETH-UCY benchmark computed based on the weighted average of a five-fold evaluation. Results are reported on $\text{minADE}_{20}$/$\text{minFDE}_{20}$. Top $\alpha\%$ split refers to performance on the highest percentile of challenging scenarios. $\text{VaR}_{99}$ and $\text{VaR}_{95}$ refer to the $0.99$ and $0.95$ quantiles of the error distribution. Relative metrics are provided in the last two columns as a normalized measure of the model performance on the long-tail. Bold numbers indicate the best performance for each metric. For all metrics lower value is better.
@@
Method Top 1 % Top 5 % All $\text{VaR}_{99}$ $\text{VaR}_{95}$ Top 1% Top 5%
Traj++ [15] 0.58/1.23 0.65/1.42 0.21/0.41 0.98/2.47 0.56/1.33 2.8/3.0 3.1/3.5
Traj++ EWTA [6] 0.45/0.88 0.44/0.91 0.18/0.32 0.75/1.92 0.47/0.92 2.5/2.8 2.4/2.8
contrastive [6] 0.40/0.71 0.44/0.91 0.19/0.32 0.74/1.82 0.47/0.93 2.1/2.2 2.3/2.8
FEND [7] 0.38/0.74 0.37/0.76 0.17/0.32 - - 2.2/2.3 2.2/2.4
AMEND(Ours) 0.27/0.44 0.25/0.41 0.14/0.22 0.63/1.53 0.36/0.67    1.9/2.0 1.8/1.9
@@ V-A Experimental Setup
[…]
@@
More formally, it is the smallest error $e$ such that the probability of observing error larger than $e$ is smaller than $1$-$\alpha$, where $E$ is the distribution of errors. We measure VaR at three levels: $0.95$, $0.97$ and $0.99$.
Baseline. We used Trajectron++ EWTA (Traj++ EWTA for short) [6] as the baseline model for our framework. Traj++ EWTA enhances Traj++ [15] by replacing its conditional Variational Auto Encoder (CVAE) module with the multi-hypothesis networks trained with Evolving Winner-Takes-All (EWTA).
Implementation Details. Our main model and the router network follow the same training schedule. We train our model for 100 epochs consisting of 20 epochs for each value of the $\textup{K}\in\{20,10,5,2,1\}$ in EWTA stages. The router softmax temperature is set to 1. The dimension of our router decoder’s hidden layer is 232. All other training details are kept the same as in the baseline model [6].
@@ V-B Experimental Variations
$p^{c}_{i}=\frac{\text{exp}(\text{-dist}(\rho(x_{i}),\phi_{c}))}{\sum_{j=1}^{C}\text{exp}(\text{-dist}(\rho(x_{i}),\phi_{j}))}$ (4)
@@
V-C Comparison to SOTA Predictors
In Table I, we compare, our proposed approach (AMEND) to several state-of-the-art pedestrian trajectory prediction models on the ETH-UCY datasets. Though our framework is designed to improve performance on the tail of the dataset, we observe that it significantly improves performance overall. This is because we effectively scaled our baseline model, increasing the number of expert models and giving each specialized abilities. Compared to the baseline, we improved performance by $22\%$ and $31\%$ on minADE and minFDE, respectively. Compared to SOTA models, we achieve the best performance on most subsets, and the best performance overall by up to $18\%$.
V-D Long-tailed Prediction
@@
From the relative metrics (the last two coloumns), we can see that given our higher performance ratio on challenging scenarios, compared to the average case, our model achieves a more balanced performance across different difficulty levels. This indicates that our approach, which consists of allocating more training resources to specific prediction patterns via specialized modules is especially helpful for complex patterns. We verify the soundness of our improvements by comparing the VaR metric which measures the error distribution quantiles. Achieving lower values on VaR metrics overall means that the largest prediction errors given by our model across the dataset is smaller than the largest errors of other models. This indicates that we have not migrated errors from one domain to another one.
V-E Training Diverse Experts
Figure 2: Radar plot showing the performance of each expert on the different cluster splits of the ETH-UCY Hotel test dataset showing significant variations in the performance of the experts. The best performing expert for each cluster tends to be the one that was assigned to it during training.
In Figure 2, we show the performance of each experts on test samples across different clusters. As expected, the experts perform best in the cluster of scenarios assigned to them during training. Note that the best performing expert for each cluster tends to be the expert that was assigned to that cluster.
In Table III, we compare different training methods to create specialized experts. For clustering (on top) we compare Trajectory, which is a model that clusters scenarios based on the final endpoints of the ego-trajectory to ours (Latent Feature). Here, we can see that our approach marginally outperforms Trajectory, especially on top $5\%$ samples. Such improvement can be due to added information captured in the latent space, accounting for factors, such as interactions between the agents.
TABLE III: The effect of clustering basis (top rows) and the routing method (bottom rows) on the overall performance and the long-tailed performance of the AMEND model. The $\text{minADE}_{20}$/$\text{minFDE}_{20}$ are reported based on the average of 5-fold evaluation for the ETH/UCY dataset. The * indicates the default setting.
Top 1 % Top 5 % All $\text{VaR}_{97}$
Trajectory 0.28/0.42 0.27/0.45 0.14/0.22 0.44/0.91
Latent Feature* 0.27/0.44 0.25/0.41 0.14/0.22 0.44/0.91
Cluster-based 0.26/0.42 0.26/0.44 0.14/0.22 0.45/1.01
Router Network* 0.27/0.44 0.25/0.41 0.14/0.22 0.44/0.91
TABLE IV: The accuracy of routing method in selecting the expert that performs best on a test sample. Best performing expert is defined as the one that achieves the lowest error metrics (ADE/FDE). Errors per sample are averaged over sixteen trials to reduce uncertainty. The * indicates our default approach. Higher value means better.
@@ V-E Training Diverse Experts
ETH - 0.32/0.37 0.35/0.36
Hotel - 0.39/0.31 0.38/0.38
Univ - 0.32/0.31 0.38/0.37
Zara1 - 0.28/0.29 0.31/0.32
Zara2 - 0.48/0.38 0.49/0.41
Avg 0.20/0.20 0.36/0.33 0.38/0.37
V-F Effectiveness of Router Network
In Table III (bottom) we compare our main model which selects experts with the router network (Router Network), to an alternative approach which selects the expert who’s sample cluster is predicted to contain the test sample (Cluster-based). Here, we can see that utilising the router network generates better results, especially on the VaR metric where improvement of up to $10\%$ is achieved.
We further investigate the discrepancy between the performance of our routing approach compared to the clustering technique. For this experiment, we report the results in terms of accuracy of the methods for predicting which expert would perform best. The results are summarized in Table IV. Here, the baseline for comparison is random routing, which has a $20\%$ chance of being correct since we have 5 experts. The strong performance of Cluster-based relative to random routing supports the hypothesis that the best performing expert for samples within a particular cluster is usually the expert specialized for the cluster. However, its lower performance compared to our Router-Network approach, suggests that there are exceptions to this rule, which our router has successfully learned. This supports the idea of using neural networks to map the complex distribution of relative expert strength across the data space for routing.
@@ VI Conclusion
In this paper, we tackled the long-tail pedestrian prediction problem by formulating a Mixture of Experts framework. We proposed a novel two-stage training scheme in which we first train specialized experts on sub-tasks within the data, and second use the experts to train a routing network for scoring the experts at inference time. We demonstrated that clustering the dataset and focusing each expert’s resources on a partition creates specialized skills which can be utilised to generate accurate predictions. We conducted extensive experimental evaluation on common pedestrian trajectory benchmark datasets and showed that our approach achieves state-of-the-art performance while outperforming the previous methods on challenging tailed samples. We further highlighted the effectiveness of our proposed modules via ablation studies.

### v2

@@ Abstract
Accurate prediction of pedestrians’ future motions is critical for intelligent driving systems. Developing models for this task requires rich datasets containing diverse sets of samples. However, the existing naturalistic trajectory prediction datasets are generally imbalanced in favor of simpler samples and lack challenging scenarios. Such a long-tail effect causes prediction models to underperform on the tail portion of the data distribution containing safety-critical scenarios. Previous methods tackle the long-tail problem using methods such as contrastive learning and class-conditioned hypernetworks. These approaches, however, are not modular and cannot be applied to many machine learning architectures. In this work, we propose a modular model-agnostic framework for trajectory prediction that leverages a specialized mixture of experts. In our approach, each expert is trained with a specialized skill with respect to a particular part of the data. To produce predictions, we utilise a router network that selects the best expert by generating relative confidence scores. We conduct experimentation on common pedestrian trajectory prediction datasets and show that our method improves performance on long-tail scenarios. We further conduct ablation studies to highlight the contribution of different proposed components.
@@ I Introduction
Our contributions are as follows: We propose a novel mixture of experts framework for trajectory prediction. Our framework encourages the diversity of expert skill sets to mitigate the long-tailed distribution problem, while simultaneously avoids additional computational cost. We conduct empirical evaluations on common pedestrian trajectory benchmark datasets and highlight the advantage of our multi-expert method on predicting challenging scenarios. At the end, we perform ablation studies, showing the benefits of proposed modules on the overall performance.
@@ IV-B Routing Samples to Experts
TABLE I: Quantitative evaluation on long-tail scenarios for the ETH-UCY benchmark computed based on the weighted average of a five-fold evaluation. Results are reported on $\text{minADE}_{20}$/$\text{minFDE}_{20}$. Top $\alpha\%$ split refers to performance on the highest percentile of challenging scenarios. $\text{VaR}_{97}$ refers to the $0.97$ quantiles of the error distribution. Relative metrics are provided in the last three columns as a normalized measure of the model performance on the long-tail. Bold numbers indicate the best performance for each metric. * indicates the model without publicly available code, hence, it is not considered in ranking. For all metrics lower value is better.
@@
Method Top 1 % Top 3 % Top 5 % All $\text{VaR}_{97}$ Top 1% Top 3% Top 5%
Traj++ 0.58/1.23 0.65/1.42 - 0.21/0.41 0.78/1.97 2.8/3.0 3.5/3.1 -
Traj++ EWTA 0.45/0.89 0.47/0.99 0.43/0.89 0.18/0.32 0.55/1.18 2.5/2.8 2.6/3.1 2.4/2.8
contrastive 0.45/0.81 0.48/0.99 0.44/0.88 0.18/0.32 0.56/1.18 2.5/2.5 2.7/3.1 2.4/2.8
FEND* 0.38/0.74 0.40/0.85 0.37/0.76 0.17/0.32 - 2.2/2.3 2.4/2.7 2.2/2.4
AMEND 0.43/0.75 0.48/0.91 0.43/0.80 0.18/0.31 0.55/1.11 2.4/2.4 2.7/2.9 2.4/2.6
@@ V-A Experimental Setup
A drawback of this metric is that the definition of the tailed scenarios is dependant on the model used to judge difficulty. Therefore, different models might underperform on different scenarios and the definition of the data tail might vary [7]. Furthermore, simply measuring errors on the set of challenging scenarios does not properly capture the changes in the distribution of errors across the dataset. It is possible for a trade-off to occur in which the model’s performance deteriorates on other scenarios. ii) Error Distribution Quantiles: We use an alternative metric which directly measures the magnitude of the tail of the error distribution. This is the error on the worst performing samples according to the model. We adopt the value-at-risk (VaR) metric. VaRα refers to the $\alpha^{th}$ quantile of the error distribution, where $\alpha\in(0,1)$:
@@
More formally, it is the smallest error $e$ such that the probability of observing error larger than $e$ is smaller than $1$-$\alpha$, where $E$ is the distribution of errors. We measure VaR at $0.97$.
Models. We report the results on two baseline models, Trajectron++ (Traj++) [15] and its variation Trajectron++ EWTA (Traj++ EWTA for short) [6], which replaces Conditional Variational Auto Encoder (CVAE) module in Traj++ with the multi-hypothesis networks trained with Evolving Winner-Takes-All (EWTA). A variation of Traj++ EWTA with contrastive loss, denoted at contrastive, is also reported. Moreover, we report on state-of-the-art model FEND [7]. However, since this model does not have a publicly available code, we do not consider it in the ranking of the models.
Implementation Details. Our main model and the router network follow the same training schedule. We train our main model for 300 epochs, with the EWTA schedule starting at $\textup{K}=20$, decaying self-adaptively by 0.8 if accuracy metrics do not improve, and ending at $\textup{K}=1$. The router softmax temperature is set to 1. The dimension of our router decoder’s hidden layer is 232. All other training details are kept the same as in the baseline model [6].
@@ V-B Experimental Variations
$p^{c}_{i}=\frac{\text{exp}(\text{-dist}(\rho(x_{i}),\phi_{c}))}{\sum_{j=1}^{C}\text{exp}(\text{-dist}(\rho(x_{i}),\phi_{j}))},$ (4)
@@
V-C Long-tailed Prediction
@@
From the relative metrics (the last two coloumns), we can see that given our higher performance ratio on challenging scenarios, compared to the average case, our model achieves a more balanced performance across different difficulty levels. This indicates that our approach, which consists of allocating more training resources to specific prediction patterns via specialized modules is especially helpful for complex patterns. We verify the soundness of our improvements by comparing the VaR metric which measures the error distribution quantiles. Achieving lower values on VaR metrics overall means that the largest prediction errors given by our model across the dataset is smaller than the largest errors of other models. For the proposed model, the improvement is apparent for final error (by 6%). This indicates that overall, our model was successful by preventing the error migration from one domain to another.
V-D Training Diverse Experts
Figure 2: Radar plot showing the FDE of each expert on the different cluster splits of the ETH-UCY Hotel test dataset showing significant variations in the performance of the experts. The best performing expert for each cluster tends to be the one that was assigned to it during training.
In Figure 2, we show the performance of each expert on test samples across different clusters. As expected, the experts perform best in the cluster of scenarios assigned to them during training. Note that the best performing expert for each cluster tends to be the expert that was assigned to that cluster.
In Table II, we compare different training methods to create specialized experts. For clustering (on top) we compare Trajectory, which is a model that clusters scenarios based on the final endpoints of the ego-trajectory to ours (Latent Feature). Here, we can see that our approach outperforms Trajectory, especially on VaR. Such improvement can be due to added information captured in the latent space, accounting for factors, such as interactions between the agents.
TABLE II: The effect of clustering basis (top rows) and the routing method (bottom rows) on the overall performance and the long-tailed performance of the AMEND model. The $\text{minADE}_{20}$/$\text{minFDE}_{20}$ are reported based on the average of 5-fold evaluation for the ETH/UCY dataset. The * indicates the default setting.
Top 1 % Top 3 % Top 5 % All $\text{VaR}_{97}$
Trajectory 0.45/0.81 0.50/0.99 0.44/0.87 0.18/0.31 0.57/1.21
Latent Feature* 0.43/0.75 0.48/0.91 0.43/0.80 0.18/0.31 0.55/1.11
Cluster-based 0.42/0.85 0.47/0.94 0.42/0.82 0.18/0.30 0.57/1.13
Router Network* 0.43/0.75 0.48/0.91 0.43/0.80 0.18/0.31 0.55/1.11
TABLE III: The accuracy of routing method in selecting the expert that performs best on a test sample. Best performing expert is defined as the one that achieves the lowest error metrics (ADE/FDE). Errors per sample are averaged over sixteen trials to reduce uncertainty. The * indicates our default approach. Higher value means better.
@@ V-D Training Diverse Experts
ETH - 0.37/0.38 0.34/0.34
Hotel - 0.40/0.36 0.40/0.39
Univ - 0.30/0.30 0.36/0.35
Zara1 - 0.34/0.34 0.34/0.35
Zara2 - 0.50/0.43 0.49/0.44
Avg 0.33/0.33 0.38/0.36 0.39/0.37
V-E Expert Selection
In Table II (bottom) we compare our main model which selects experts with the router network (Router Network), to an alternative approach which selects the expert who’s sample cluster is predicted to contain the test sample (Cluster-based). Here, we can see that utilising the router network generally generates better results.
We further investigate the discrepancy between the performance of our routing approach compared to the clustering technique. For this experiment, we report the results in terms of accuracy of the methods for predicting which expert would perform best. The results are summarized in Table III. Here, the baseline for comparison is random routing, which has a $33\%$ chance of being correct since we have 3 experts. The strong performance of Cluster-based relative to random routing supports the hypothesis that the best performing expert for samples within a particular cluster is usually the expert specialized for the cluster. However, its lower performance compared to our Router-Network approach, suggests that there are exceptions to this rule, which our router has successfully learned. This supports the idea of using neural networks to map the complex distribution of relative expert strength across the data space for routing.
@@ VI Conclusion
In this paper, we tackled the long-tail pedestrian prediction problem by formulating a Mixture of Experts framework. We proposed a novel two-stage training scheme in which we first train specialized experts on sub-tasks within the data, and second use the experts to train a routing network for scoring the experts at inference time. We demonstrated that clustering the dataset and focusing each expert’s resources on a partition creates specialized skills which can be utilised to generate accurate predictions. We conducted extensive experimental evaluation on common pedestrian trajectory benchmark datasets, outperforming the previous methods on challenging tailed samples. We further highlighted the effectiveness of our proposed modules via ablation studies.